[Halla_running] [EXTERNAL] Hall A RC Daily Summary August 19th, 2025
McMurtry, Jacob Thomas (rby2vw)
rby2vw at virginia.edu
Tue Aug 19 09:01:48 EDT 2025
Dear Collaborators,
Over the last 24 hours we have continued to take production data on the LH2 target with an uptime of 79%. ECal and HCal experts lead by Andrew implemented new gain coefficients for the ECal. Afterwards, they did the same for the HCal. Two sets of runs at 10 uA, one with 25% HCal threshold and one with 40% threshold, were taken. We are now running with 70% ECal and 40% HCal thresholds (https://logbooks.jlab.org/entry/4445518 ). GEM experts continued their tests on the two modules (https://logbooks.jlab.org/entry/4445413 ), but could not run through the night due to gas issues (https://logbooks.jlab.org/entry/4445688 ). During the OWL, the TO noticed a dip in the target scattering chamber pressure and called the target expert (https://logbooks.jlab.org/entry/4445774 ). At this time it does not appear to be a concern, but we should keep an eye out for any further drops. Accelerator had some downtime incidents related to the storm last night: https://logbooks.jlab.org/entry/4445431 , https://logbooks.jlab.org/entry/4445777 , https://logbooks.jlab.org/entry/4445792 . The current run plan is to continue taking production data at 15 uA on LH2.
MCC is doing RF recovery tomorrow. Most recent estimate is that it's expected to take 4-6 hours, and probably won't open up the accelerator for any work. Halls B and D have requested rapid and controlled access for this time. There are currently three requests from safety and Moller folks for access.
Since yesterday's Day to today's owl:
* Total Accumulated Charge at Kin3 so far: 86.94 C (Goal: 100 C)
* Total Accumulated Charge at Kin3 since yesterday : 0.83 C
Updates from MCC:
* Tomorrow is RF recovery day
* Compton test is scheduled for Thursday day shift (estimated 8 hours, no other beam studies planned)
* The Energy Secretary tour is this Thursday, he will be touring Hall C in the morning and Hall B in the afternoon (current schedule)
* No update or news on high power test
Current Issues and Hall activities:
* ECal/HCal calibration
* Status of GEM gas tests
Shift Summary:
* Yesterday's Day: https://logbooks.jlab.org/entry/4445104
* Yesterday's Swing: https://logbooks.jlab.org/entry/4445643
* Today's Owl: https://logbooks.jlab.org/entry/4445646
RC Meeting:
* We will have our usual RC meeting today at 1 pm in the conference room in the second floor of the CH and on Zoom.
* https://urldefense.proofpoint.com/v2/url?u=https-3A__jlab-2Dorg.zoomgov.com_j_1616220064&d=DwIFAw&c=CJqEzB1piLOyyvZjb8YUQw&r=cE_5H9I241Ivk1xpqB8Fq0UiLTkf4zMIQ6sedLYb2m8&m=2JHW7oIjhFcvWZqjq8_4n6AxWemslaxWyxRFwedpld2OznI9OX6odmjZXiXlfNrn&s=2aXtTI3lgU1V-FrdjpzOYj783clPYHA6MPVDFccwlx4&e=
* Meeting ID: 161 622 0064
* Passcode: 314159
Best,
Jacob McMurtry
Hall A RC
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halla_running/attachments/20250819/7916b154/attachment.htm>
More information about the Halla_running
mailing list